What is the probability of drawing a yellow marble or a red marble from a bag containing 12 yellow marbles, 16 red marbles, and 15 green marbles?
Answer by checkley71(8403)

THESE TYPE OF PROBABILITIES ARE SIMPLY [SUCCESS COUNT/TOTAL CHANCES]
(12+16)/(12+16+15)=28/43=.65 OR 65% CHANCE OF DRAWING A YELLOW OR RED MARBLE.